OpenStudy (anonymous):

find the inverse of f(x)=(2/3)x-(1/4)

OpenStudy (anonymous):

finding the inverse of the function is simply switching the variable x and y and solve the equation in terms of y.. let f(x)=y \(y= \frac{2x}{3}- \frac{1}{4}\) switch x and y it will be \(x= \frac{2y}{3}- \frac{1}{4}\) then solve this new equation in terms of y..